Use the double angle identity to find sin(2x) if csc= -square root 19 and cosx>0
Answer by Alan3354(69443)

Use the double angle identity to find sin(2x) if csc= -square root 19 and cosx>0
--------
sin = 1/csc = -1/sqrt(19)
---
cos = sqrt(1 - sin^2) = sqrt(18)/sqrt(19)
---
sin(2x) = 2sin(x)cos(x)
= 2*(-1)*sqrt(18)/19
= -6sqrt(2)/19